UT November 19 2018

OSA: J. Williams

ISA: B. Rothberg (Tucson)

AO Support (FLAO): G. Taylor (Tucson)

AO Support (ARGOS): S. Walsh

Laser Dude: G. Rahmer

Software Support: A. Colson

Mtn Manager: R. Hanson

ISp: G. Bechetti

UT 06:55 - reconfigured for LUCI-LUCI/ARGOS Bino - pointing check

Test dark exposures taken with both LUCIs (0001-0002)

UT 06:59 - clouds to the north and southwest on all-sky, they were NOT there when LN handed over

UT 07:02 - w-unit on DX was powered off - Greg trying to recover

UT 07:03 - sending colliation bino script for Palomar 2

PCS issue seems to have resolved for tonight so far

UT 07:08 - clouds thickening....

UT 07:13 - sent ARGOS laser preset

UT 07:14 - mountain dropped out again

UT 07:16 - paused - trying to close loops - also satellite

UT 07:17 - pointing through thin clouds - double star resolved on CCD47, DX AdSec shell RIP'd

UT 07:24 - still having issue - gain changes and shell rips ... no data taken yet.

UT 07:28 - thicker layer of clouds - wiping out AO Ref star and affecting the lasers

UT 07:29 - switching to try ESM on same object

UT 07:35 - sending bino preset - SL left ESM right

UT 07:37 - DX missed the guidestar - doing a pointing check

After re-authorize to Bino-LUCI pointing was way off on DX

UT 07:47 - sending bino preset SL-SX ESM-DX

UT 07:52 - Gustavo is turning on the lasers.....because....

UT 07:57 - resending preset - AdSec had a failure on a step in the acquisition process

UT 07:59 - dichroic did not move out - ARGOS lockout shows on GUI - but dichroic out shows on IIF

UT 08:00 - something odd going on with DX AdSec - Greg checking

UT 08:02 - we won't be running lasers at same time....since the light blows up CCD47....this is what dichroic protects from

UT 08:04 - restarting DX AO

UT 08:08 - resending preset

UT 08:09 - preset on right failed - didn't find guidestar

UT 08:13 - all-sky shows clear - going back to ARGOS


UT 08:15 - authorizing ARGOS on DX then doing a pointing check

UT 08:33 - sending ARGOS collimation preset

UT 08:35 - turning lasers on - failed to faind GS on left (SX)

UT 08:41 - sending collimation preset

UT 08:45 -missed on SX - no GS found

UT 08:51 - sending collimaiton preset again

UT 08:53 - turning on lasers

UT 08:57 - sending science preset BINO SX-SL DX-ARGOS

SX WFS shows 0.5"

Double star on CCD47 is a single blob - seems to be screwing up collimation - ESM was OK with it 2 nights ago

L1 0003-0004 OK

L2 0003-0004 bad exposures - show poor collimation

UT 09:03 - sending revised script with a new AO Ref star (slightly fainter R=14)

UT 09:08 - sending revised scripts

UT 09:16 - revised a 4th time - have to hand edit and hand adjust pointing center

L1 0005-0008 OK data (but never made it to an off field

L2 0005 - looks bad - low gain

L2 0006 - Looks OK

L2 0007 - AdSeC failure

L2 0008 - bad

UT 09:23 - inserted a pause in script

UT 09:29 - Lost AORef star on DX - resending preset

-preset failed on DX but we see AORef star on CCD47

UT 09:31 - sending preset again

L1 0009-0020

L2 0009-0020

UT 09:38 Co-pointing Error Again with big offset

UT 09:40 - disabling co-pointing check - sending offset

Syncoffset received, but no feedback,

UT 09:42 - telescope service shows "Abort Wait for Preset" - clicked it

IIF Shows "timeoutException - timeout while sending receiving data

UT 09:44 - Manually sending offset with Telescope Service - waiting for flight to pass

UT 09:48 - sent offset via telsvc - clicked continue on LUCI -script

WORKAROUND CO-POINTING ERROR with MIXED MODE:- fails with co-pointing disabled, need to manually put in the offsets via Telsvc then commit - continue LUCI script from where it stopped and GO

UT 09:53 - script commanding us to move back to science field - went through OK

L2 0017-0018 stars jumped -

L2 0019 - bad frame - stars jumped

L2 0020 - DX AdSec RIP'd

UT 09:57 - paused script before offset - when fixed will repeat the on-source dither

UT 10:01 - additional errors on DX AdSec - DD issue as well

UT 10:02 - SX WFS 0.51"

UT 10:03 - lost AORef Star on DX - sending preset again

L1 0021 (new preset to Pal2) -

L2 0021 (new preset to Pal2) -

UT 10:05 - SX WFS 0.48"

UT 10:06 - DX AdSec RIP'd - inserted pause

L2 0020-0021 - bad

UT 10:12 - restarting exposures from first dither position (0,0)

Image quality on L2 look elongated at 15-30 degree angle (NE)

UT 10:16 offset to sky frames worked

L2 - 0027 jumpd - but this is a sky-frame so not a big deal

UT 10:18 - laser paused for airplane (already paused before for satellite)

UT 10:19 - contining in sky frame

UT 10:21 - need to reacquire AO Ref star

L2 0031-0032 bad - will have to redo the on-field images

UT 10:23 - DX AdSec RIP'd, inserted pause - if need to resend preset will restart at dither position -45 -30

UT 10:25 - Shane working on laser issue

UT 10:27 - AOS crashed

UT 10:29 - AdSec recovered - resending preset

NOTE FOR ARGOS: Scripts made need to insert pauses without taking data - to help with recovery or acquisition

SX WFS 0.46"

UT 10:32 - skipping to -45 30 offset in science field

L2 0033-0034 - stars jumped

L2 0035-0036 - stars jumped

L2 0037 - first off shows jumped too

L2 0038 - 2nd off shows jump too

UT 10:42 - had to skip to line to force the return to field

L2 0041 - images elongated

All L2 images have elongation regardless of on/off science field - PSFGUI right shows tip/tilt being applied

UT 10:56 - reconfigured for LUCI-BINO SL - running SL Bino script to test (since L1 is OK, this will constrain if this is an L2/DX issue or related to ARGOS - if SL is OK, then we will try ESM with L2)

UT 10:57- missed the guidestar on the right (DX)

UT 10:58 - pointing check

UT 11:07 - sending SL Bino preset to Palomar2

UT 11:15 - SL L2 looks fine (round ~0.5" FWHM)

L1 0051-0062 SL Kband test

L2 0051-0062 SL Kband test

- made new script for ESM that is same field center as ARGOS and SL with same exact offset

UT 11:17 - called Greg to let him know we will be running ESM

UT 11:25 - sent SL/ESM script

UT 11:41 - DX AO arbitrator was wonky - resending preset

UT 11:52 - DX AdSec RIP'd

UT 12:26 - need to set shells - moiving to higher elevation

UT 12:35 - Palomar 2 getting low, AdSec fixed, moving to M67 for SL/ESM test - pointing check

UT 12:38 - sending m67 preset

UT 12:40 - complained that GS was out of limits on SX (but it was happy on DX)

UT 12:44 - same complaint - new GS is withing ARGOS-GS field so no way its out of limits - Josh is homing AGW1 and I'll resend

UT 12:46 - sending again

UT 12:48 - processes are being run as ARGOS - could this be causing problems while running ESM?

UT 12:49 - sending M67 yet again

UT 12:51 - authorizing as ARGOS LUCI becuase swithing to FLAO will take too long with -12 coming

UT 12:55 - another complaint about GS in ARGOS patrol field - homing AGW1 (2 was happy) and now resending with different star

UT 12:56 - sending with new ARGOS-GS

UT 12:56 - dichroic failed to move on LUCI-2

UT 12:58 - IIF thinks the dichroic is out - showing up as an error

UT 12:59 - reauthorizing for ARGOS-LUCI-2 - to see if this clears error

UT 01:06 - Wolfgang and Gustavo trying to figure out error with dichoric (BCU issue?)

UT 01:08 -pointing on right side off

UT 01:10 - pointing check to fix pointing issue

UT 01:17 - going to another pointing star - can't get star on DX-side

UT 01:27 - pointing OK, but ARGOS issues not fixed - would take too long - finishing up

UT 01:28 - closing up

L1 0063-

L2 0063-

-- BarryRothberg - 19 Nov 2018
Topic revision: r1 - 19 Nov 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback